OpenStudy (anonymous):

evaluate each expression: Logs. Click 2 see.

OpenStudy (anonymous):

\[\log _{\sqrt{8}}4096\]

OpenStudy (anonymous):

is that sqrt{8} or {3}

OpenStudy (anonymous):

8

OpenStudy (anonymous):

\[\huge \log _{\sqrt{8}}{4096}=\log_{8^{1/2} }4096\] use change of base rule\[\huge \log _ba=\frac{\log a}{\log b}\]

OpenStudy (anonymous):

okay . and then simplify that ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay then?

OpenStudy (anonymous):

you mean take 4096 and divide it by 8^1/2 ?

OpenStudy (anonymous):

yes but both preceded by log

OpenStudy (anonymous):

4096=64^2 \[\huge \frac{\log 4096}{\log 8^{1/2}}\]

OpenStudy (anonymous):

i got 4 ?

OpenStudy (anonymous):

@Jonask

OpenStudy (anonymous):

\[\huge \frac{4\log 8}{\log 8}=4\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

but the back of my book has the answer as: 8 ?

OpenStudy (anonymous):

i forgot the half also it is supposed to be dropped\[\huge \frac{4 \log 8}{1/2\log 8}=4\div{1/2}=8\]

OpenStudy (anonymous):

OH okay . thanks :)

OpenStudy (anonymous):

yw